SUPERSIMPLE ω-CATEGORICAL GROUPS

FRANK O WAGNER

Abstract. An ω-categorical supersimple group is finite-by-abelian-by-finite, and has finite SU-rank. Every definable subgroup is commensurable with an acl(∅)- definable subgroup.